# Almost Surely Safe Alignment of Large Language Models at Inference-Time

### v2

@@ 7 Conclusion
We introduced InferenceGuard, a novel inference-time alignment method that ensures large LLMs generate safe responses almost surely—i.e., with a probability approaching one. We extended prior safety-augmented MDP theorems into the latent space of LLMs and conducted a new analysis. Our results demonstrated that InferenceGuard significantly outperforms existing test-time alignment methods, achieving state-of-the-art safety versus reward tradeoff results. In the future, we plan to improve the algorithm’s efficiency further and generalize our setting to cover jailbreaking. While our method is, in principle, extendable to jailbreaking settings, we aim to analyze whether our theoretical guarantees still hold.
Broader Impact Statement
Acknowledgments
The authors would like to thank Rasul Tutunov, Abbas Shimary, Filip Vlcek, Victor Prokhorov, Alexandre Maraval, Aivar Sootla, Yaodong Yang, Antonio Filieri, Juliusz Ziomek, and Zafeirios Fountas for their help in improving the manuscript. We also would like to thank the original team from Pekin University, who made their code available and ensured we could reproduce their results.
@@ Appendix A Additional Related Work
Safe RL: Safe RL employs the cMDP framework (Altman, 1999) to enforce safety constraints during exploration and policy optimization. When no prior knowledge is available, methods focus on safe exploration (Turchetta et al., 2016; Koller et al., 2018; Dalal et al., 2018; Wachi & Sui, 2018; Bharadhwaj et al., 2020). With prior knowledge, such as environmental data or an initial safe policy, methods learn safe policies using control techniques like Lyapunov stability (Chow et al., 2018; 2019; Berkenkamp et al., 2017; Ohnishi et al., 2019) and reachability analysis (Cheng et al., 2019; Akametalu et al., 2014; Dean et al., 2019; Fisac et al., 2019). Safety constraints are enforced via Lagrangian or constrained optimization methods (Achiam et al., 2017; Ray et al., 2019; Stooke et al., 2020; Yang et al., 2019; Ding et al., 2020; Ji et al., 2024b), but can often lead to suboptimal safety-reward trade-offs. In contrast, our approach extends safety state augmentation Sootla et al. (2022) to LLMs and latent MDPs to ensure almost sure inference time safety without relying on Lagrangian multipliers.
@@
LLM alignment and safety: Methods for aligning pre-trained LLMs with task-specific data include prompting, guided decoding, and fine-tuning. Among fine-tuning methods, RL from Human Feedback (RLHF) has proven effective, where LLMs are fine-tuned with a learned reward model (Stiennon et al., 2020; Ziegler et al., 2019; Ouyang et al., 2022) or directly optimized from human preferences (Rafailov et al., 2023; Azar et al., 2023; Zhao et al., 2023; Tang et al., 2024; Song et al., 2024; Ethayarajh et al., 2024; Ramesh et al., 2024). Recent works have explored fine-tuning for helpful and harmless responses (Bai et al., 2022; Ganguli et al., 2022), while (Dai et al., 2023) introduced a safe RL approach incorporating safety cost functions via Lagrangian optimization, requiring model weight fine-tuning. Other safety-focused methods, including machine unlearning (Gundavarapu et al., 2024), safety pre-aligned multi-modal LLMs (Gou et al., 2024), safety-aware model merging (Hammoud et al., 2024), prompting-based safety methodologies (Hua et al., 2024), test-time controllable safety alignment (Zhang et al., 2024b), defenses against adversarial attacks and jailbreaking (Guo et al., 2024; Qi et al., 2024; Xu et al., 2024), identifying safety criticial regions in LLMs (Wei et al., 2024), safety preserved LoRA fine-tuning (Li et al., 2025a), alignment using correctional residuals between preferred and dispreferred answers using a small model (Ji et al., 2024a), and identifying safety directions in embedding space (Feng et al., 2024).Those methods are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety during inference.
@@
Inference time alignment: The closest literature to ours is inference-time alignment. Those methods offer flexible alternatives to fine-tuning LLMs, as they avoid modifying the model weights. A common approach is guided decoding, which steers token generation based on a reward model. In particular, (Khanov et al., 2024; Shi et al., 2024; Huang et al., 2024) perform this guided decoding through scores from the reward model whereas Han et al. (2024); Mudgal et al. (2023); Kong et al. (2024) use a value function that is trained on the given reward model. These inference-time alignment methods build on previous works like (Yang & Klein, 2021; Arora et al., 2022; Krause et al., 2021; Kim et al., 2023; Meng et al., 2022; Peng et al., 2019), which guide or constrain LLMs towards specific objectives. Other safety-focused inference-time methods include, reverse prompt contrastive decoding (Zhong et al., 2024b), adjusting model hidden states combined with guided decoding (Banerjee et al., 2024), soft prompt-tuned detoxifier based decoding (Niu et al., 2024), jail-break value decoding (Wang et al., 2024a), speculating decoding using safety classifier (Zeng et al., 2024), and opposite prompt-based contrastive decoding (Zhao et al., 2024). Compared to those methods, we are the first to achieve almost sure safe alignment with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints (see Section 5) without introducing additional parameters like Lagrangian multipliers.
@@ Appendix B Theoretical Analysis
For our theoretical results, we consider a similar setup to that of Sootla et al. (2022); Hernández-Lerma & Muñoz de Ozak (1992) but with a discrete action space. Consider an MDP $M=\{S,A,P,c,\gamma_{c}\}$ with a discrete, non-empty, and finite action set for each $s$ defined as $A(s)$. The set
$K=\{(s,a)\mid s\in S,a\in A(s)\}$
@@
Note that, the assumptions B1-B3, share a similar essence to that of the Assumptions 2.1-2.3 in Hernández-Lerma & Muñoz de Ozak (1992) and B1-B3 in Sootla et al. (2022) but suited for a discrete action space. In particular, Assumption B3, is similar to the lower semi continuity assumption on the set-value function map $A(s)$ taken in Sootla et al. (2022); Hernández-Lerma & Muñoz de Ozak (1992) but modified for a discrete action space.
Our first goal is to recreate Hernández-Lerma & Muñoz de Ozak (1992, Lemma 2.7) for our discrete action setting. Let $\Pi$ denote the set of functions from $S\to A$.
@@
  • (a)
    If Assumption B3 holds and $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a\in A(s)$ and bounded from below on $K$, then the function
@@
  • (b)
    If the Assumptions B1-B3 hold, and $u\in L(S)$ is nonnegative, then the (nonnegative) function
@@
  • (c)
    For each $n=0,1,\dots$, let $v_{n}$ be a l.s.c. function, bounded from below. If $v_{n}\to v_{0}$ as $n\to\infty$, then
@@
For part a)
We have $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a$. This implies from the definition of lower semi -continuity for any $s_{0}\in S$ and $a\in A(s_{0})$, if $v(s_{0},a)>y$, then there exists a $\epsilon>0$, s.t. $\forall s$ satisfying $\|s-s_{0}\|\leq\epsilon$, $v(s,a)>y$.
@@
This proves the lower semi continuity of $v^{*}(s)=\inf_{a\in A(s)}v(s,a)$. Further, due to the discrete nature of $A(s)$, the $\inf_{a\in A(s)}v(s,a)$ is always attained by an action $\pi(s)\in A(s)$. Hence, there exists a function $\Pi:S->A(s)$ s.t.
@@
For part b), note that $c(s,a)+\int u(y)P(dy\mid s,a)$ is l.s.c. w.r.t. $s$ for an given $a$, based on Assumptions B1-B2. Hence, using part a) we have,
$u^{*}(s):=\inf_{a\in A}\left[c(s,a)+\int_{S}u(y)P(dy\mid s,a)\right]=c(s,\pi(s))+\int_{S}u(y)P(dy\mid s,\pi(s))\in L(S)\quad\forall s\in S$ (14)
@@ B.1 Optimality Equation
Next, we characterize the solution to the Bellman (Optimality) equation. We begin by recalling the Bellman operator:
@@
Note that, by induction and Lemma 1(b) again, $v_{h}\in L(S)^{+}$ for all $h\geq 0$. From elementary Dynamic Programming Bertsekas (1987); Bertsekas & Shreve (1996); Dynkin & Yushkevich (1979), $v_{h}(s)$ is the optimal cost function for an $h$-stage problem (with ”terminal cost” $v_{0}(s)=0$) given $s_{0}=s$; i.e.,
@@
We want to prove similar results to that of Hernández-Lerma & Muñoz de Ozak (1992, Theorem 4.2) on the optimality of the Bellman operator, however in the discrete action setting. In particular, we want to show the following theorem
@@
We follow a similar proof strategy to that of Hernández-Lerma & Muñoz de Ozak (1992, Theorem 4.2).
@@
where $P^{H}(B\mid s,\pi)=P(\{s_{H}\in B\})$ denotes the $H$-step transition probability of the Markov chain $\{s_{h}\}$ (see Hernández-Lerma & Muñoz de Ozak (1992, Remarks 3.1,3.2)). Therefore, since $u$ is nonnegative,
@@
This implies $u(s)\leq V^{*}(s)$. We have thus shown that $u=V^{*}$.
Further, if there is another solution $u^{\prime}$ satisfying $u^{\prime}=Tu^{\prime}$, it holds that $u^{\prime}\geq V^{*}$, Hence, $V^{*}$ is the minimal solution.
@@ B.2 Limit of a sequence of MDPs
Consider now a sequence of Markov Decision Processes (MDPs) $M_{n}=\{S,A,P,c_{n},\gamma_{n}\}$, where, without loss of generality, we write $c_{n},c_{\infty}$ and $M_{n},M_{\infty}$. Consider now a sequence of value functions $\{V_{n}^{*}\}_{n=0}^{\infty}$:
$V_{n}(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{n}(s_{t},a_{t})\right],$
$V_{n}^{*}(s)=\inf_{\pi}V_{n}(\pi,s).$
The ”limit” value functions (with $n=\infty$) are still denoted as follows:
$V(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],$
$V^{*}(s)=\inf_{\pi}V(\pi,s).$
@@
In addition to the previous assumptions, we make an additional one, while modifying Assumption B1:
  • B1’
    For each $n$, the functions $c_{n}(s,a)$ are bounded, measurable on $\mathcal{K}$, nonnegative, lower semi-continuous;
  • B4
    The sequence $\{c_{n}(s,a)\}_{n=0}^{\infty}$ is such that $c_{n}\uparrow c$.
@@
We follow a similar proof strategy to that of Hernández-Lerma & Muñoz de Ozak (1992, Theorem 5.1)
@@ B.3 Latent MDP Analysis
We begin by comparing our assumptions to that of the assumptions B1-B4, closely aligned to those used in Hernández-Lerma & Muñoz de Ozak (1992); Sootla et al. (2022).
To prove a),b) of Definition 4.1 we need to verify that the latent MDP satisfying Assumptions A1-A2 also satisfies Assumptions B1’, B2-B4. According to Assumption A1, we consider bounded costs $\bar{\mathcal{C}}^{n}_{\text{task}}$ continuous w.r.t. state $(\mathbf{h},\mathbf{o},\text{z})$ for a given $y$ with discrete and finite action space ${\mathcal{V}}$, hence Assumptions B1’, B3, and B4 are satisfied. Assumptions B2 and A2 are identical. This proves a) and b).
@@
$V(\bar{\pi}(\phi(\cdot)),\mathbf{s}_{0},\text{z}_{0})$ (33)
$=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t+1},\text{z}_{t+1}\sim{\mathcal{P}}(\cdot|\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\tilde{\mathcal{C}}^{n}_{\text{task}}(\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\Big{]}$ (34)
@@ Appendix C Algorithmic details
Input: Initial prompt $s_{0}$, beam depth $d$, $N$ number of beams, max depth $D$, top $K$ beams, $M$ max retry
@@
for $\frac{D}{K}$ iterations do
      $\forall i\in\{1,\dots d\},j\in\{1,...,|\mathcal{V}|\},F_{i,j}\leftarrow$ 0

### v3

@@ 7 Conclusion
We introduced InferenceGuard, a novel inference-time alignment method that aims to ensure LLMs generate safe responses almost surely w.r.t. given safety cost model. We extended prior safety-augmented MDP theorems into the latent space of LLMs and conducted a new analysis. Our results demonstrated that InferenceGuard significantly outperforms existing test-time alignment methods, achieving state-of-the-art safety versus reward tradeoff results.
Limitations. Our almost sure safety guarantees rely on the quality of the given safety cost model and, hence, improving its safety assessment ability remains a subject for future investigation. We also plan to improve the algorithm’s efficiency further and generalize our setting to cover jailbreaking. While our method is, in principle, extendable to jailbreaking settings, we aim to analyze whether our theoretical guarantees still hold.
@@ Appendix A Additional Related Work
Safe RL: Safe RL employs the cMDP framework [72] to enforce safety constraints during exploration and policy optimization. When no prior knowledge is available, methods focus on safe exploration [73, 74, 75, 76, 77]. With prior knowledge, such as environmental data or an initial safe policy, methods learn safe policies using control techniques like Lyapunov stability [78, 79, 80, 81] and reachability analysis [82, 83, 84, 85]. Safety constraints are enforced via Lagrangian or constrained optimization methods [65, 86, 87, 88, 89, 70], but can often lead to suboptimal safety-reward trade-offs. In contrast, our approach extends safety state augmentation [16] to LLMs and latent MDPs to ensure almost sure inference time safety without relying on Lagrangian multipliers.
@@
LLM alignment and safety: Methods for aligning pre-trained LLMs with task-specific data include prompting, guided decoding, and fine-tuning. Among fine-tuning methods, RL from Human Feedback (RLHF) has proven effective, where LLMs are fine-tuned with a learned reward model [2, 17, 11] or directly optimized from human preferences [21, 22, 23, 24, 25, 26, 90]. Recent works have explored fine-tuning for helpful and harmless responses [27, 8], while [28] introduced a safe RL approach incorporating safety cost functions via Lagrangian optimization, requiring model weight fine-tuning. Other safety-focused methods, including machine unlearning [29], DPO with expanded safety zone [91], dual objective DPO with targeted unlearning [92], safety pre-aligned multi-modal LLMs [30], safety-aware model merging,editing [31, 93, 94], prompting-based safety methodologies [32, 95, 96], test-time controllable safety alignment [33], defenses against adversarial attacks and jailbreaking [34, 97, 35, 98, 99, 100, 101, 102, 103], identifying safety critical regions in LLMs [36], safety preserved LoRA fine-tuning [37], alignment using correctional residuals between preferred and dispreferred answers using a small model [104], layer-specific editing of LLMs to ensure safety [105], training a small model to correct the outputs of a large model [106], and identifying safety directions in embedding space [107]. Those methods are either orthogonal, handle a different problem to ours, or can not ensure almost sure safety during inference.
@@
Inference time alignment: The closest literature to ours is inference-time alignment. Those methods offer flexible alternatives to fine-tuning LLMs, as they avoid modifying the model weights. A common approach is guided decoding, which steers token generation based on a reward model. In particular, [43, 44, 45] perform this guided decoding through scores from the reward model whereas Han et al. [46], Mudgal et al. [14], Kong et al. [42] use a value function that is trained on the given reward model. These inference-time alignment methods build on previous works like [40, 108, 109, 110, 111, 112], which guide or constrain LLMs towards specific objectives. Other safety-focused inference-time methods include, reverse prompt contrastive decoding [47], adjusting model hidden states combined with guided decoding [113, 114], cross-model guidance through safety steering vectors [115], adjusting logits based on self-evaluation [116, 35], soft prompt-tuned detoxifier based decoding [49], jailbreak-targeted inference-time interventions[53, 117], speculating decoding using safety classifier [50], comparing the cosine similarity between the target prompt and a set of harmful and benign prompts to classify for refusal [118], decoding using two fine-tuned models (one trained for safety and other as an adversary) [119, 120], and opposite prompt-based contrastive decoding [48]. Compared to those methods, we are the first to achieve almost sure safe alignment with strong empirical results. Operating in the latent space enables us to train smaller, inference-efficient critics while optimally balancing rewards and safety constraints (see Section 6) without introducing additional parameters like Lagrangian multipliers.
@@ Appendix B Theoretical Analysis
For our theoretical results, we consider a similar setup to that of Sootla et al. [16], Hernández-Lerma and Muñoz de Ozak [15] but with a discrete action space. Consider an MDP $M=\{S,A,P,c,\gamma_{c}\}$ with a discrete, non-empty, and finite action set for each $s$ defined as $A(s)$. The set
$K=\{(s,a)\mid s\in S,a\in A(s)\}$ (10)
@@
Note that, the assumptions B1-B3, share a similar essence to that of the Assumptions 2.1-2.3 in Hernández-Lerma and Muñoz de Ozak [15] and B1-B3 in Sootla et al. [16] but suited for a discrete action space. In particular, Assumption B3, is similar to the lower semi continuity assumption on the set-value function map $A(s)$ taken in Sootla et al. [16], Hernández-Lerma and Muñoz de Ozak [15] but modified for a discrete action space.
Our first goal is to recreate Hernández-Lerma and Muñoz de Ozak [15, Lemma 2.7] for our discrete action setting. Let $\Pi$ denote the set of functions from $S\to A$.
@@
  • (a) If Assumption B3 holds and $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a\in A(s)$ and bounded from below on the set $K$ (see Equation 10), then the function
@@
  • (b) If the Assumptions B1-B3 hold, and $u\in L(S)$ is nonnegative, then the (nonnegative) function
@@
  • (c) For each $n=0,1,\dots$, let $v_{n}$ be a l.s.c. function, bounded from below. If $v_{n}\to v_{0}$ as $n\to\infty$, then
@@
For part a, note that, we have $v(s,a)$ is l.s.c. w.r.t. $s$ for any given $a$. This implies from the definition of lower semi-continuity for any $s_{0}\in S$ and $a\in A(s_{0})$, if $v(s_{0},a)>y$, then there exists a $\epsilon>0$, s.t. $\forall s$ satisfying $\|s-s_{0}\|\leq\epsilon$, $v(s,a)>y$.
@@
This proves the lower semi continuity of $v^{*}(s)=\inf_{a\in A(s)}v(s,a)$. Further, due to the discrete nature of $A(s)$, the $\inf_{a\in A(s)}v(s,a)$ is always attained by an action $\pi(s)\in A(s)$. Hence, there exists a function $\Pi:S->A(s)$, such that,
@@
For part b), note that $c(s,a)+\int u(y)P(dy\mid s,a)$ is l.s.c. w.r.t. $s$ for an given $a$, based on Assumptions B1-B2. Hence, using part a) we have $\forall s\in S$,
$u^{*}(s):=\inf_{a\in A}\left[c(s,a)+\int_{S}u(y)P(dy\mid s,a)\right]=c(s,\pi(s))+\int_{S}u(y)P(dy\mid s,\pi(s))\in L(S)$ (15)
@@ B.1 Optimality Equation
In this section, we characterize the solution to the Bellman (Optimality) equation. We begin by recalling the Bellman operator:
@@
Note that, by induction and Lemma 1(b) again, $v_{h}\in L(S)^{+}$ for all $h\geq 0$. From elementary Dynamic Programming [121, 122, 123], $v_{h}(s)$ is the optimal cost function for an $h$-stage problem (with "terminal cost" $v_{0}(s)=0$) given $s_{0}=s$; i.e.,
@@
We want to prove similar results to that of Hernández-Lerma and Muñoz de Ozak [15, Theorem 4.2] on the optimality of the Bellman operator, however in the discrete action setting. In particular, we want to show the following theorem
@@
We follow a similar proof strategy to that of Hernández-Lerma and Muñoz de Ozak [15, Theorem 4.2].
@@
where $P^{H}(B\mid s,\pi)=P(\{s_{H}\in B\})$ denotes the $H$-step transition probability of the Markov chain $\{s_{h}\}$ (see Hernández-Lerma and Muñoz de Ozak [15, Remarks 3.1,3.2]). Therefore, since $u$ is nonnegative,
@@
This implies $u(s)\leq V^{*}(s)$. We have thus shown that $u=V^{*}$. Further, if there is another solution $u^{\prime}$ satisfying $u^{\prime}=Tu^{\prime}$, it holds that $u^{\prime}\geq V^{*}$, Hence, $V^{*}$ is the minimal solution. ∎
@@ B.2 Limit of a sequence of MDPs
Consider a sequence of Markov Decision Processes (MDPs) $M_{n}=\{S,A,P,c_{n},\gamma_{n}\}$, where, without loss of generality, we write $c_{n},c_{\infty}$ and $M_{n},M_{\infty}$ with corresponding value functions $\{V_{n}^{*}\}_{n=0}^{\infty}$:
$V_{n}(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c_{n}(s_{t},a_{t})\right],\quad V_{n}^{*}(s)=\inf_{\pi}V_{n}(\pi,s).$
The "limit" value functions (with $n=\infty$) are still denoted as follows:
$V(\pi,s_{0})=\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right],\quad V^{*}(s)=\inf_{\pi}V(\pi,s).$
@@
In addition to the previous assumptions, we make an additional one, and modify Assumption B1:
B1’ For each $n$, the functions $c_{n}(s,a)$ are bounded, measurable on $\mathcal{K}$, nonnegative, lower semi-continuous w.r.t. $s$ for a given $a\in A(s)$;
B4 The sequence of cost functions $\{c_{n}(s,a)\}_{n=0}^{\infty}$ converges to $c$, i.e, $c_{n}\uparrow c$.
@@ B.3 Latent MDP Analysis
We begin by comparing our assumptions to that of the assumptions B1-B4, closely aligned to those used in Hernández-Lerma and Muñoz de Ozak [15], Sootla et al. [16].
To prove a),b) of Definition 5.1 we need to verify that the latent MDP satisfying Assumptions A1-A2 also satisfies Assumptions B1’, B2-B4. According to Assumption A1, we consider bounded costs $\bar{\mathcal{C}}^{n}_{\text{task}}$ continuous w.r.t. state $(\mathbf{h},\mathbf{o},\text{z})$ for a given $y$ with discrete and finite action space ${\mathcal{V}}$, hence Assumptions B1’, B3, and B4 are satisfied. Assumptions B2 and A2 are identical. This proves a), b).
@@
$V(\bar{\pi}(\phi(\cdot)),\mathbf{s}_{0},\text{z}_{0})$ $=\mathbb{E}_{\begin{subarray}{c}\mathbf{s}_{t+1},\text{z}_{t+1}\sim{\mathcal{P}}(\cdot|\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\\ \text{y}_{t}\sim\bar{\pi}(\cdot|\phi(\mathbf{s}_{t}),\text{z}_{t})\end{subarray}}\Big{[}\sum_{t=0}^{\infty}\gamma^{t}\tilde{\mathcal{C}}^{n}_{\text{task}}(\mathbf{s}_{t},\text{z}_{t},\text{y}_{t})\Big{]}$ (33)
@@ Appendix C Algorithmic details
Input: Initial prompt $s_{0}$, beam depth $d$, $N$ number of beams, max depth $D$, top $K$ beams, $M$ max retry, reference policy $\pi_{\text{ref}}$
@@
for $\frac{D}{d}$ iterations do
      $\forall i\in\{1,\dots d\},j\in\{1,...,|\mathcal{V}|\},F_{i,j}\leftarrow$ 0, $\pi_{\text{pen}}\leftarrow\pi_{\text{ref}}$
@@
Here, $\pi_{\text{pen}}$ is the penalized policy that for each step $t+i\in[t,t+d]$, samples from $\text{SoftMax}(\mathbf{W}\mathbf{o}_{t+i}-\text{n}_{2}({F_{i}>0}))$ where $\text{n}_{2}({F_{i}>0})$ is a vector where each component $j$ is $\text{n}_{2}$ if $F_{j,i}>0$ and 0 otherwise. This avoids sampling the same token at the same position observed in unsuccessful beams, thus increasing diversity.